OpenStudy (anonymous):

help please.... anyone know how to work this out: The circle C, with centre at the point A, has equation x2 + y2 – 10x + 9 = 0. Find (a) the coordinates of A, - i tried subsituting in x=0 but it didnt work?

OpenStudy (anonymous):

** \[x^2+y^2-10x+9=0\]

OpenStudy (anonymous):

\[(x-a)^2+(y-b)^2=r^2\] \[x^2-10x+y^2+9=0\]\[x^2-10x+y^2+9+16-16=0\]\[x^2-10x+y^2+25=4^2\]\[(x-5)^2+(y-0)^2=4^2\] here we can see the coordinates for A in the y direction will be 0 and for the x direction will be 5 So A is at the point (5,0)

OpenStudy (anonymous):

where did you get the 16 from?

OpenStudy (anonymous):

we have to complete the square to get it in the right form

OpenStudy (anonymous):

@kinzan123, That method is called "Completing the Square".

OpenStudy (anonymous):

thanks :) ...with completing the square, you start by factorising right?

OpenStudy (anonymous):

yes but it's of a special form (x-a)^2 that's why you complete the square

OpenStudy (anonymous):

yeh.. so you get (x+4)^2 right?

OpenStudy (anonymous):

then you have to takewaway the 16 to make it equal?

OpenStudy (anonymous):

close so we had x^2-10x+9 let's ignore the y for now,, x^2-10x+9+16-16 =x^2-10x+25-16 =(x-5)^2-16 because (x-a)^2=x^2-2ax+a^2

OpenStudy (anonymous):

No, Completing the Square actually completes the square (just as what it is called). Factorization comes at a later point when you have completed the square. The process involves a binomial of a variable plus a constant (not always permanently). Check this example. If you have \[(x+4)^{2}\] That would then be equal to \[x^2+8x+16\] Reversing the process would be Completing the Square. Taking for example \[x^2−10x+y^2+9=0\] You add 16 in order to make 9 be 25, and then subtract 16 in order to balance the equation \[x^2−10x+y^2+9+16−16=0\] At this point, you can then have 9+16 as 25 and transpose (sorry about this term) 16 to the right side of the equation and simplify it to its square root squared. \[x^2−10x+y^2+25=4^2\] Then group the equation \[x^2−10x+25+y^2=4^2\] Then factorization comes along: \[(x−5)^2+y^2=4^2\] Then just simplify further and you have it. Hope you understood.

OpenStudy (anonymous):

@sercczionelabus but why would you want to make 25? is it because you factoeised the x terms to get (x-5)^2

OpenStudy (anonymous):

Because of the middle term that is -10. If you have \[ax^2+bx+c\] Use this to get the constant that would make the quadratic equation a perfect squared binomial \[(b/2)^2\] Let's take a look back at the previous equation. \[x^2−10x+y^2+9=0\] Ignoring y and using the formula I gave you, you would have \[(-10/2)^2=25\] So you must have a constant of magnitude 25 in order to complete the square, and since you've got 9 already; all you have to do then is add 16.

OpenStudy (anonymous):

ohhh , so in general ..do i always use (b/2)^2 ??

OpenStudy (anonymous):

Yea. It is a very useful formula. Though it's not a standard formula, it is just a formula based on analysis.

OpenStudy (anonymous):

oh okay, thank you very much :)

OpenStudy (anonymous):

Sure, no problem.

OpenStudy (anonymous):

how would we simplify the following further?: \[(x-5)^2+y^2=4^2\]

OpenStudy (anonymous):

That is actually very much simplified. But since you're dealing with a circle, you have to take note of the general form, \[(x-h)^2+(y-k)^2\] (Not sure if I have interchanged k and h). So you can then add 0 to the left side of the equation (wouldn't matter since it is a zero and would still render the equation balanced). And\[y^2+0\] is just the same as\[(y+0)^2\] Upon writing it that way, you have now formulated it in such a way that it matches the general form. Then from h and k, you can then take the coordinates of your A (I guess, not sure about this since I forgot about this).

OpenStudy (anonymous):

I forgot to put =r^2 at the last part of the general form. Sorry, my bad.

OpenStudy (anonymous):

lol thanks :)

OpenStudy (anonymous):

Sure.